OpenStudy (anonymous):

find all real solutions for sin^2x-2sinx=0

OpenStudy (anonymous):

Factor out a power of sine: \[\sin^2x-2\sin x=0~~\iff~~\sin x(\sin x-2)=0\]

OpenStudy (anonymous):

thank you!